Doubling time of an investment with continuous compound interest is 12.4 years. If the investment is worth $23000, how much will it be worth 5 years from now?
Answer by lwsshak3(11628)

Doubling time of an investment with continuous compound interest is 12.4 years. If the investment is worth $23000, how much will it be worth 5 years from now?
***
compound interest formula for continuous compounding
A=Pe^rt, P=initial investment, r=interest rate, A=amt after t-years
A/P=e^rt
A/P=2
2=e^rt
2=e^r*12.4
take log of both sides
ln2=12.4r*lne
0.693=12.4r
r=.056
..
5-yrs from now:
A=23000*e^.056*5
A=30432
how much will $23000 be worth 5 years from now? $30432
***